HIGH RESOLUTION FOURIER TRANSFORM FAR-INFRARED SPECTROSCOPY OF CH_{3}OD: GLOBAL FIT OF TORSION-ROTATIONAL TRANSITIONS IN THE FIRST THREE TORSIONAL STATES

\underline{INDRANATH MUKHOPADHYAY}, Laser Programme, Centre for Advanced Technology, Indore 452 013, India; XIAN-XIAO HAO, Department of Physics, Yantai Normal University, Yantai, Shandong 264000, China; GEORG CHR. MELLAU and STEFAN KLEE, Physikalisch Chemisches Institut der Justus-Liebig-Universitt, Heinrich-Buff-Ring 58, D-35392 Giessen, Germany; YUN-BO DUAN and ANNE B. MCCOY, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

The results of a new analysis of the high resolution Fourier transform far infrared (FIR) absorption spectrum of the torsionrotational band of CH_{3}OD will be presented. Based on a recently determined set of molecular constants for CH_{3}OD^{a} we have been able to assign 1126 assigned Fourier transform far-infrared (FIR) transitions that involve the second excited torsional levels with v_{t} = 2. A CH_{3}OD data set that contains 460 microwave (MW), millimeter wave (MMW) transitions and 3474 Fourier transform FIR transitions with v_{t} ≤ 2 and J ≤ 21 has been fit using a reduced torsion-rotational Hamiltonian obtained from the one-large-amplitude internal rotation model. The MW and MMW transitions have been fit with a root-mean-square (rms) deviation of 0.12 MHz while FIR transitions have a rms deviation of 0.00026 cm^{−1} using 61 parameters. These deviations are on the order of the experimental uncertainties, indicating that the MW, MMW and FIR spectral transitions have been fit to the desired accuracy and the reduced torsion-rotational Hamiltonian model is capable of accurately describing CH_{3}OD energy levels up through the second excited torsional level. The success of the fit demonstrates that the increased general asymmetry in CH_{3}OD, compared to CH_{3}OH, can be accounted for adequately by the reduced Hamiltonian model.

aY. B. Duan and A. B. McCoy, J. Mol. Spectrosc., 199, 302-306 (2000).